MEASUREMENT OF ROTATIONAL STATE-TO-STATE RELAXATION COEFFICIENTS BY RAMAN-RAMAN DOUBLE RESONANCE. APPLICATION TO SELF-COLLISIONS IN ACETYLENE.

\underline{J. L. DOMENECH}, R. Z. MARTINEZ, and D. BERMEJO, Instituto de Estructura de la Materia (CSIC), Serrano 123, E-28006 Madrid, SPAIN.

We have developed a technique for the measurement of state-to-state rotational relaxation rates due to collisions in the gas phase. A single J state of the v_{2} = 1 vibrational level of acetylene is populated by a Stimulated Raman process. After a variable delay of a few ns we record high resolution (∼ 0.003 cm^{−1}) spectra of the Q-branch of the v_{2} = 2 ← v_{2} = 1 transition by a second Stimulated Raman process. The relative intensities of the lines of this Q-branch, recorded as a function of the the initially pumped J level and the number of collisions between the pump and probe stages (delay × pressure product), allows us to obtain a set of state-to-state energy transfer coefficients by fitting the evolution of the observed populations to that predicted by a master equation with adjustable coefficients. The experimental details and first results for odd-J levels of acetylene at 150 K will be presented.